40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 174 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  8340   Mon Mar 25 18:07:31 2013 SteveUpdate40m UpgradingETMY table is out

Quote:

Quote:

Yes! We are swapping.

I'll be there very soon!

 The ETMY optical table 4' x 2' with all optical components was placed on two carts and rolled out of the end.  The 4' x 3' x 4" TMC 78-235-02R breadboard was placed into position and marked for anchoring bolt locations.

It will be drilled, tapped, shimmed, leveled and bolted it into final position tomorrow morning.  I'm planning to bring  the acrylic enclosure to the east end tomorrow afternoon.

 The new table is tapped and leveled, but not ready for final anchoring. The existing 8 mm shims on the top of the support legs will be moved to the bottom of the legs, where more torque is available on the 1/2" bolts

The coated tin film sheets are being installed at the shop. The table and the enclosure will be ready on Friday.

Attachment 1: shimmingEE.jpg
shimmingEE.jpg
  8339   Mon Mar 25 16:51:59 2013 JenneUpdateOpticsupdated calculations of PRC/SRC g-factors and ARM mode matching

I think we like the idea of flipping SR2 better than SR3 for the same ghost beam reasons as PR2 is better than PR3.  There isn't very much free space in the BS chamber, and if we flip PR/SR3, we have to deal with green ghosts as well as IR.

The flipping SR2 case seems okay to me - flipping either one of the SR folding mirrors gives us a slightly better g-factor than the design with infinite curvatures, and flipping SR2 gives us slightly better mode matching to the arm than the flipped SR3 case, but more importantly, there are fewer ghosts to deal with.  I vote we flip SR2, not SR3.

  8338   Mon Mar 25 16:51:43 2013 ChloeUpdate Final QPD Circuit Design

This is the final version of the QPD circuit I'm going to build. After playing around with the spatial arrangement, this should fit into the box that I was planning to use, although it will be a rather tight fit. The pitch, yaw, and summing circuit will be handled with a quad op amp. Planning to meet with Eric tomorrow to figure out the logistics of building things.

In the meantime, I'm reading about designing the ECDL for my summer project with Tara. He sent me several papers to read so we can talk on Wednesday.

Attachment 1: IMG-20130325-00244.jpg
IMG-20130325-00244.jpg
  8337   Mon Mar 25 15:24:41 2013 ranaUpdateLockingTT1 drift

 

 I doubt that it has truly disappeared. Can you please make a measurement of it and quantify the hysteresis using some angle sensor?

  8336   Mon Mar 25 15:01:33 2013 ManasaUpdateLockingTT1 drift

Quote:

2. TT1 drifting in pitch (Bistable)
During the arm alignment routine for spot centering, we observed that TRY dropped (from TRY = 0.9 until the arm lost lock) every 40minutes or so. The arm was relocked by moving TT1 in pitch. The (locking - unlocking due to drift - relocking) cycle was monitored and we observed that it was bistable i.e. if TT1 was moved up in pitch (0.2 on the slider) to relock for the first time ; the next time it lost lock, TT1 had to be moved down by nearly the same distance to relock the arm.
Moving TT2 or the testmasses did not help with relocking the arms; so TT1 seems to be the one causing all th trouble atleast for today.

Update

TT1 was moved in pitch to restore flashing in the arms on Wednesday (Mar 20) so that it doesn't drift too far off making it difficult to lock again. Since then, the arms have been flashing without any drifting and TTs have remained untouched. The hysteresis has disappeared.

  8335   Mon Mar 25 11:42:45 2013 JamieUpdateComputersc1lsc mx_stream ok

I'm not exactly sure what the problem was here, but I think it had to do with a stuck mx_stream process that wasn't being killed properly.  I manually killed the process and it seemed to come up fine after that.  The regular restart mechanisms should work now.

No idea what caused the process to hang in the first place, although I know the newer RCG (2.6) is supposed to address some of these mx_stream issues.

  8334   Mon Mar 25 09:52:22 2013 JenneUpdateComputersc1lsc mxstream won't restart

Most of the front ends' mx streams weren't running, so I did the old mxstreamrestart on all machines (see elog 6574....the dmesg on c1lsc right now, at the top, has similar messages).  Usually this mxstream restart works flawlessly, but today c1lsc isn't working.  Usually to the right side of the terminal window I get an [ok] when things work.  For the lsc machine today, I get [!!] instead. 

After having learned from recent lessons, I'm waiting to hear from Jamie.

  8333   Fri Mar 22 23:23:38 2013 AnnalisaUpdateLockingBeat note still missing

[Annalisa, Manasa, Koji]

I updated the setup for the beat note. The main reason is that I needed to keep the ADJ to 0 in way to operate at the nominal laser power.

Now the input power of the laser is increased (about 315 mW) and needs to be dumped so as not to exceed the PD threshold of 1mW. 

Moreover, a lens has been added to match the two beams size.

A BS has been removed from the PSL pick-off beam path, so the PSL power hitting the BS is now about 100 uW, and the total power on the PD is 0.7mW.

I also verified that both the beams are S polarized.

To find the beat note, the laser temperature has been varied  through the laser controller and not adding a Voltage with the power supply.

A range of temperature of 30 degC has been spanned, but we suppose there should be some calibration problem with the controller, since set temperature is not the same as Laser temperature on the display. 

Anyway, no beat note has been found up to now.

 

An external monitor has to be added to check the real temperature of the crystal.

The next possible plan is to vary the PSL temperature and try to find the beat note. 

 

P.S.: The HEWELETT PACKARD 8591E spectrum analyzer works! The monitor only took some time to turn on!

 

Attachment 1: Beat_note_-_new_setup.jpg
Beat_note_-_new_setup.jpg
  8332   Fri Mar 22 19:46:29 2013 KojiSummaryLSCDiode impedance test result

I've tested Perkin-Elmer InGaAs PDs at OMC Lab.

- The diode impedances were measured with the impedance measurement kit. Reverse bias of 5V was used.

- Diode characteristics were measured between 10MHz and 100MHz.

- 4-digit numbers are SN marked on the can

- Ls and Rs are the series inductance and resistance

- Cd is the junction capacitance.

- i.e. Series LCR circuit o--[Cd]--[Ls]--[Rs]--o

C30665GH, Ls ~ 1nH

0782 Perkin-Elmer, Rs=8.3Ohm, Cd=219.9pF
1139 Perkin-Elmer, Rs=9.9Ohm, Cd=214.3pF
0793 Perkin-Elmer, Rs=8.5Ohm, Cd=212.8pF

C30642G, Ls ~ 12nH

2484 EG&G, Rs=12.0Ohm, Cd=99.1pF
2487 EG&G, Rs=14.2Ohm, Cd=109.1pF
2475 EG&G glass crack, Rs=13.5Ohm, Cd=91.6pF
6367 ?, Rs=9.99Ohm, Cd=134.7pF
1559 Perkin-Elmer, Rs=8.37Ohm, Cd=94.5pF
1564 Perkin-Elmer, Rs=7.73Ohm, Cd=94.5pF
1565 Perkin-Elmer, Rs=8.22Ohm, Cd=95.6pF
1566 Perkin-Elmer, Rs=8.25Ohm, Cd=94.9pF
1568 Perkin-Elmer, Rs=7.83Ohm, Cd=94.9pF
1575 Perkin-Elmer, Rs=8.32Ohm, Cd=100.5pF

C30641GH, Perkin Elmer, Ls ~ 12nH

8983 Perkin-Elmer, Rs=8.19Ohm, Cd=25.8pF
8984 Perkin-Elmer, Rs=8.39Ohm, Cd=25.7pF
8985 Perkin-Elmer, Rs=8.60Ohm, Cd=25.2pF
8996 Perkin-Elmer, Rs=8.02Ohm, Cd=25.7pF
8997 Perkin-Elmer, Rs=8.35Ohm, Cd=25.8pF
8998 Perkin-Elmer, Rs=7.89Ohm, Cd=25.5pF
9000 Perkin-Elmer, Rs=8.17Ohm, Cd=25.7pF

 

Note:  Calculated Ls&Rs of straight wires
  1mm Au wire with dia. 10um -> 1nH, 0.3 Ohm
20mm BeCu wire with dia. 460um -> 18nH, 0.01 Ohm

  8331   Fri Mar 22 01:28:56 2013 ManasaUpdateLasersBeam profile of NPRO from ATF

The NPRO from ATF has been installed on the POY table.

I have been making measurements to characterize the beam profile of this laser. I am using an AR coated laser window as a beam sampler at 45deg and the razor blade technique to measure the beam size along z. Details of the procedure along with analysis and results from this will follow.

  8330   Thu Mar 21 17:06:34 2013 SteveUpdate40m UpgradingETMY table is out

Quote:

Yes! We are swapping.

I'll be there very soon!

 The ETMY optical table 4' x 2' with all optical components was placed on two carts and rolled out of the end.  The 4' x 3' x 4" TMC 78-235-02R breadboard was placed into position and marked for anchoring bolt locations.

It will be drilled, tapped, shimmed, leveled and bolted it into final position tomorrow morning.  I'm planning to bring  the acrylic enclosure to the east end tomorrow afternoon.

Attachment 1: ETMYtableSupport.jpg
ETMYtableSupport.jpg
  8329   Thu Mar 21 14:02:45 2013 AnnalisaUpdateLockingincident angles

Quote:

Is there a reason to use non-45 degree incident angles on the steering mirrors between the laser and the PD?  I would always use 45 degree incident angles unless there is a really good reason not to.

 Actually not, it is a mistake! It is one of the things I'm going to modify, in addition to add more BS to reduce the power. 

  8328   Thu Mar 21 13:50:08 2013 JamieUpdateLockingincident angles

Is there a reason to use non-45 degree incident angles on the steering mirrors between the laser and the PD?  I would always use 45 degree incident angles unless there is a really good reason not to.

  8327   Thu Mar 21 13:11:42 2013 AnnalisaUpdateLockingFinding the beat note

Quote:

Give us more info on the elog:
What PD are you using? How much power the beams on the recombining BS are? What kind of BS is it?
How are you looking for the beat note? (on the scope? or spectrum analyzer?)
What was the scanned temp range?

Three points to be checked:

- Polarization

- Alignment

- Temperature

Experimental Setup

I'm using a 1611 New Focus PD (1 GHZ, with maximum input power 1mW), and the total power hitting on the PD is of about 0.650 mW.

The current of the NPRO laser is set to 1.38 A, so that the input power is 19 mW. The beam is initially damped by a 10% reflection BS and then it hits a 33% reflection BS (where it recombines with the PSL pick-off beam) with 2 mW power.

After this second BS the power is reduced to 0.592 mW.

The PSL pick-off hits on the 33% reflection BS with 65.5 uW power, and it exit with a 47 uW power.

I connected a power supply to apply a Voltage to the slow frequency BNC, in way to tune the laser frequency.

I'm using the AGILENT 4395A Spectrum analyzer to make the measurement. I tried to use the HEWELETT PACKARD 8591E spectrum analyzer, but the monitor didn't turn on.

The temperature spanned until now in only of about 10 deg C, because I realized that I needed a better alignment, so I added a lens in front of the PD and I did a better alignment. 

Moreover, the current of the laser is too low, so I need to increase it and add more beam splitters in the beam path to dump the beam, in way to don't reach the PD threshold.

I knew that both the beams are s-polarized, but maybe I can check it again.

Attachment 1: Beat_note_setup.jpg
Beat_note_setup.jpg
  8326   Thu Mar 21 12:33:51 2013 ranaUpdateComputersFixed

Please stop power cycling computers. This is not an acceptable operation (as Jamie already wrote before). When you don't know what to do besides power cycling the computer, just stop and do something else or call someone who knows more. Every time you kill the power to a computer you are taking a chance on damaging it or corrupting some hard drive.

In this case, the right thing to do would be to hook up the external keyboard and monitor directly to c1sus to diagnose things.

NO MORE TOUCHING THE POWER BUTTON.

  8325   Thu Mar 21 12:04:05 2013 ManasaUpdateComputersFixed

All FE computers are back.

Restart procedure:

0a. Restart frame builder: telnet fb 8087 & type shutdown

0b. Restart mx_stream from the FE overview screen

1. I ssh ed to the computer. (c1lsc, c1ioo, c1iscex, c1isey)

2. I used 'sudo shutdown -r (computername)'. They came back ON.

3. While rebooting c1ioo, c1sus shutdown (for reasons I don't know). I could not ping or ssh c1SUS after this.

4. I went in and switched c1SUS computer OFF and back ON after which I could ssh to it.

5. I did the same reboot procedure for c1SUS.

6. I had to restart some of the models individually.
    (i) ssh to the computer running the model
    (ii) rtcds restart 'model name'

7. All computers are back now.

up.png

  8324   Thu Mar 21 10:29:12 2013 ManasaUpdateComputersComputers down since last night

I'm trying to figure out what went wrong last night. But the morning status...the computers are down.

 

down.png

Attachment 1: down.png
down.png
  8323   Thu Mar 21 10:19:28 2013 KojiUpdateLockingFinding the beat note

Give us more info on the elog:
What PD are you using? How much power the beams on the recombining BS are? What kind of BS is it?
How are you looking for the beat note? (on the scope? or spectrum analyzer?)
What was the scanned temp range?

Three points to be checked:

- Polarization

- Alignment

- Temperature

  8322   Thu Mar 21 09:53:46 2013 AnnalisaUpdateLockingFinding the beat note

 Yesterday I tried to find the beat note between the main PSL and the auxiliary NPRO, but I didn't :( 

Today I will do a better alignment of the two beams in the PD and try again.

  8321   Thu Mar 21 08:46:46 2013 ManasaUpdate40m UpgradingETMY table swap

Yes! We are swapping.

I'll be there very soon!

  8320   Thu Mar 21 08:11:53 2013 SteveUpdate40m UpgradingETMY table swap

Quote:

Quote:

As discussed at the 40m meeting: Koji, Manasa and Steve

We are planning to remove the whole 4'x2' optical table ETMY-ISCT with optics as it is tomorrow morning.
This way I can start placing the new 4'x3' table and acrylic enclosure in place.

I will be removing all cables on the ETMY endtable and labeling them today before we remove the table tomorrow morning. If there is anything else that should be done before the swapping which we might have not considered, elog it and we'll have it all done.

Also,I've attached the updated inventory.

 The cables are Not disconnected. I have people coming to do the lifting at 9:45am  Email is not working. This will cause a delay. I need conformation that we still moving on with the swap

  8319   Wed Mar 20 16:45:59 2013 ManasaBureaucracyAuxiliary lockingFetched NPRO from ATF

[Koji, Annalisa, Manasa]

NPRO with controller from ATF joins the 40m. We have put it on the POY table where we plan to use it for ABSL.

Attachment 1: P3203298.JPG
P3203298.JPG
Attachment 2: P3203299.JPG
P3203299.JPG
  8318   Wed Mar 20 15:51:54 2013 ManasaUpdate40m UpgradingEndtable upgrade

Quote:

As discussed at the 40m meeting: Koji, Manasa and Steve

We are planning to remove the whole 4'x2' optical table ETMY-ISCT with optics as it is tomorrow morning.
This way I can start placing the new 4'x3' table and acrylic enclosure in place.

I will be removing all cables on the ETMY endtable and labeling them today before we remove the table tomorrow morning. If there is anything else that should be done before the swapping which we might have not considered, elog it and we'll have it all done.

Also,I've attached the updated inventory.

Attachment 1: Endtable_inventory.pdf
Endtable_inventory.pdf
  8317   Wed Mar 20 14:52:09 2013 ManasaUpdateSAFETYMC trans RFPD missing

PDA255 at the MC transmission end is missing and whoever had removed it earlier has left the beam undamped.

It could be less than a mW beam...but it is still not acceptable to leave it that way.

On the other note, I have cleared up the unwanted optics from the same table.

MC.png

  8316   Wed Mar 20 14:44:47 2013 JamieUpdateOpticsupdated calculations of PRC/SRC g-factors and ARM mode matching

Below are new alamode calculations of the PRC and SRC g-factors and arm mode matchings.  These include fixes to the ABCD matrices for the flipped folding mirrors that properly (hopefully) take into account the focusing effect of passing through the optic substrates.

I've used nominal curvatures of -600 m for the G&H PR2/SR2 optics, and -700 m for the Laseroptik PR3/SR3 dichroics.

An interesting and slightly disappointing note is that it looks like it actually would have been better to flip PR3 instead of PR2, although the difference isn't too big.  We should considering flipping SR3 instead or SR2 when dealing with the SRC.  I'll take responsibility for messing up the calculation for the flipped TTs.

PRC

PR2 Reff PR3 Reff PRC g-factor (t/s) ARM mode matching
Inf Inf .94/.94 .999
-600 -700 .99/.98 .84
413 (flipped) -700 .94/.93 .999
-600 409 (flipped) .92/.94 .999
413 (flipped) 409 (flipped) .87/.89 .97

SRC

SR2 Reff SR3 Reff SRC g-factor (t/s) ARM mode matching
Inf Inf .96/.96 .999
-600 -700 NA NA
413 (flipped) -700 .96/.95 .998
-600 391 (flipped) .94/.96 .996
413 (flipped) 391 (flipped) .90/.92 .96

RXA: maybe nominal, but we don't actually have measurements of the installed optics' curvatures, so there could be ~10-15% errors in the RoC. Which translates into a 1-2% error in the g-factor.

  8315   Wed Mar 20 14:00:21 2013 SteveUpdate40m UpgradingEndtable upgrade

As discussed at the 40m meeting: Koji, Manasa and Steve

We are planning to remove the whole 4'x2' optical table ETMY-ISCT with optics as it is tomorrow morning.
This way I can start placing the new 4'x3' table and acrylic enclosure in place.

  8314   Wed Mar 20 08:59:24 2013 SteveUpdatePEMair cond maitenance is today

 Outside air quality at 8:30am  1.1 million of 0.5 micron particles cf / min with MET counter #2

AC unit output measured  directly at the outlet of east end  6,000  and south end 3,000 of 0.5 micron particles cf / min with counter #2

Lab measurement at the top of IOOC 20K of 0.5 micron particles cf / min with MET #1

HEPA filters: 2 mobile units  at the middle of the east arm at speed 100%, PSL enclosure 20% and south end flow bench are running.

The PSL, south flow bench and mobile unit (Envirco  sn69406003 ) measured zero counts of 0.5 & 0.3 micron particles

Envirco HEPA filter unit sn 69406001 measured 50 particles of 0.5micron and 320 of 0.3 micron particles cf / min with counter #2

This filter will be replaced.

                   

NOTE: inside air quality peaks we can not control. It depends on the weather and our house keeping.

When your work is particle sensitive keep your eyes on the particle counter.

 

 

 

 

Attachment 1: 100dPEM.png
100dPEM.png
  8313   Tue Mar 19 20:24:56 2013 AnnalisaUpdateAuxiliary lockingAuxiliary laser on PSL table

 I moved the auxiliary laser from the ITMY table to the PSL table and installed all the optics (mirrors and lenses) to steer the beam up to a PDA55 photodiode, where also the pick-off of the PSL is sent.

Tomorrow I'm going to measure the beat note between the two.

  8312   Tue Mar 19 19:19:50 2013 JenneUpdate40m UpgradingEndtable upgrade : INVENTORY

Does "Ready (in stock)" for the base plates mean in stock at the company, or in hand and already here in the lab?

For the 2" mirror mounts, are we planning on making a decision soon, or a long time in the future....i.e. is this something that should go in the temporary solution column?

For the mode matching and collimating lenses, do we actually have the ones you want in our lens kits?  A lot of those kits are half empty.

  8311   Tue Mar 19 17:00:14 2013 SteveUpdateVAClarge window quote of Cascade

Optical quality 6.5" OD window specs and quotes for vacuum view ports and optical table enclosure are on the 40m wiki_ Aux Optics page.

We have just received a very good quote #2  from Cascade  Optical.

  8310   Tue Mar 19 14:21:57 2013 ManasaUpdate40m UpgradingEndtable upgrade : INVENTORY

THE INVENTORY FOR ETMY ENDTABLE.

We definitely need the polaris mirror mounts and cylindrical lenses. As for the rest, we can still go ahead and do the swapping without them.

 

Attachment 1: Endtable_inventory.pdf
Endtable_inventory.pdf
  8309   Tue Mar 19 10:30:00 2013 steveUpdateSAFETYcrane inspection 2013

 

Professional crane inspector: Fred Goodbar found two small leaks at the Vertex trolley as he was conducting the annual inspection of the 40m cranes.  Otherwise the cranes are in safe condition.

Attachment 1: sippingOIL.jpg
sippingOIL.jpg
Attachment 2: 03191301.PDF
03191301.PDF
  8308   Mon Mar 18 20:13:18 2013 AnnalisaBureaucracyAuxiliary lockingYend table upgrade - go fetch NPRO from ATF

Quote:

1) Annalisa is going to start  working on mode profiling and beat note search for the old MOPA NPRO.

2) In the meantime, Manasa is working on the end table items. This will be reviewed by KA in the afternoon.

The laser at ATF is moved to the 40m when the status of 1) and 2) is determined by KA to be reasonable.

We also make the beat note measurement for the ATF laser too.
 

Today I installed mirrors to steer the pick-off from the main laser beam in a more free part of the PSL table and make the beat note measurement between it and the NPRO.

At the beginning I took the beam from the harmonic separator after the doubling crystal, and I was going to bring it in a less full part of the table . At the end I realized that there was already a beam steered up to a more free part of the table, and the beam is taken from the transmission of the PMC.

Tomorrow I'm going to use that beam to find the beat note with the NPRO.

I also removed almost all the steering  optics that I used on the ITMY table to send the auxiliary beam for ABSL through the window parallel to the POY beam. The most important thing is that I removed the BS, which was on the same path of the POY beam (see elog 8257).

 

  8307   Mon Mar 18 14:58:15 2013 SteveUpdateGreen Lockingrepair or replace dead NPRO

 

 JDSU can repair the Lightwave M126-1064-700 NPRO, sn 415  They do not need the Controller sn 516 

 Posted in the 40m Wiki_ PSL_ NPRO  cost repair and/or option to buy Innolight laser as replacement

  8306   Mon Mar 18 13:10:19 2013 KojiBureaucracyAuxiliary lockingYend table upgrade - go fetch NPRO from ATF

1) Annalisa is going to start  working on mode profiling and beat note search for the old MOPA NPRO.

2) In the meantime, Manasa is working on the end table items. This will be reviewed by KA in the afternoon.

The laser at ATF is moved to the 40m when the status of 1) and 2) is determined by KA to be reasonable.

We also make the beat note measurement for the ATF laser too.
 

  8305   Mon Mar 18 12:35:29 2013 AnnalisaBureaucracyAuxiliary lockingYend table upgrade - go fetch NPRO from ATF

Quote:

Zach has just replied, and said that we should feel free to take the laser from his iodine setup in the West Bridge subbasement, in the ATF lab. 

Annalisa, please ask Koji or Tara to show you where it is, and help you bring it to the 40m.  You should install it (temporarily) on the PSL table, measure the waist, and find the beat in IR.  Elog 3755 and elog 3759 have some of the details on how it has been done in the past.

 Ok, I'm going to contact Koji.

  8304   Mon Mar 18 12:23:25 2013 JenneBureaucracyAuxiliary lockingYend table upgrade - go fetch NPRO from ATF

Zach has just replied, and said that we should feel free to take the laser from his iodine setup in the West Bridge subbasement, in the ATF lab. 

Annalisa, please ask Koji or Tara to show you where it is, and help you bring it to the 40m.  You should install it (temporarily) on the PSL table, measure the waist, and find the beat in IR.  Elog 3755 and elog 3759 have some of the details on how it has been done in the past.

  8303   Mon Mar 18 12:02:12 2013 AnnalisaConfigurationABSLABSL setup for g-factor measurement of PRC
Motivations
The ABSL technique has been already used in the past to measure the absolute length of the interferometer's optical cavities by means of an auxiliary laser source, as described in LIGO-P1200048-v3 and in Alberto Stochino thesis work.
Using the same technique it is possible to measure the g-factor of the power recycling cavity by measuring the cavity Transverse Mode Spacing.
 
Plan for experimental setup
The auxiliary laser is set on the POY table and is injected through the ITMY window in way to follow the same path of the POY beam. It hits the AR wedge of ITMY and is reflected back to the BS and the PRM.
 
Since the main beam is P-polarized, all the optics in the central IFO are P-polarization dependent, so it is useful to P-polarize the auxiliary beam before it enters the IFO.  
I made a mode matching calculation with a la mode script, in order to mode match the auxiliary beam waist to the waist of the main laser.
However, before ordering and installing steering optics and mode maching lenses, I'm waiting to know whether someone has an NPRO laser to install on the END table in place of the broken one, otherwise the one I'm using could be taken.
In this case a possibility could be to take the auxiliary beam from the end table with an optical fiber, but it means to use the auxiliary laser alternately to lock the arm or make a measurement of TMS. If so, a new calculation for the mode matching needs to be done.
Anyway, I hope that another laser will be found!
 
In order to phase lock the auxiliary beam with the main beam, the latter will be taken from the PSL table after the PMC through a single mode fiber, which will be brought up to the POY table. This solution results to be more reliable then taking the POY beam to phase lock the two laser, because POY is related to the locking. 
 
The signal with the beat note between the two lasers can be detected by the transmission from PR2 (POP). 
 
 
 
  8302   Fri Mar 15 16:46:52 2013 JenneBureaucracyAuxiliary lockingYend table upgrade - fast track?

In light of the Yend auxiliary laser's ill health, I think we should reconsider the possibility of changing out the Yend laser table next week.

My thinking here is that if whatever the new mode matching solution is for a replacement laser (Tara has borrowed our spare NPRO that used to sit on top of the fridge, or we could take Annalisa's) requires a rework of the table layout, we might as well put the new layout onto the new table.  So, we need to figure out what laser we will put in as the new Ygreen, and what it's waist looks like.  If it just requires a small movement of existing lenses or new lenses in similar positions to the current ones, we can keep living with our current table.  But, if the mode matching solution requires enough changes to distances / lens placement / whatever, we should think seriously about putting in the new table next week.

Here's what I would like to see happen on / before Monday:

Annalisa - Mode matching solution for new laser.  If we get the laser back from Tara, this will involve first measuring the waist, otherwise we already know the waist of the ABSL laser that Annalisa is currently using.

Annalisa and Steve - Find optics for new mode matching in the lab, or order them by Monday afternoon.

Manasa - List of every screw, washer, optic, mount, etc. that will go on the new Y end table, with a notation as to whether or not we have it in-hand, and if not, what needs to happen before we do.  Also, for things that we don't have, I'd like to see a summary of temporary solutions (e.g. keep using current mount for doubling crystal while new one is being machined).

Manasa / Annalisa / Koji - will the new mode matching solution fit within the existing layout, or do we need to redo the table layout?

  8301   Fri Mar 15 15:26:13 2013 KojiUpdateGreen LockingYarm ALS laser is funny / dying

I took a look at the laser. It is probably the LD TEC (DTEC) failure.
As the temperature of the LD (DTMP) gradually deviated from 25degCish,
the DTEC voltage also went up from 2Vish to 2.1, 2.2... 

When DTEC reaches 3V, it stopped lasing. This cools the diode a bit, and
it start lasing but repeat the above process.

I am not sure which of the head and controller has the issue.

The situation did not improve much by reducing the pumping current (ADJ: -15).

BTW, Turning on/off the noise eater did not change the situation.

I think the head/controller set should be sent out to JDSU and find how they will say.

  8300   Fri Mar 15 02:19:01 2013 JenneUpdateGreen LockingYarm ALS laser is funny / dying

I turned the laser back on around 1am.  This is still happening, although right now it is turning off more often than before, maybe every 15 seconds or so.  I am going to turn off the laser for the night.

The measured laser temperature is about 45C (I have a 25,000 count offset in the Y ALS Slow control right now....higher offset, lower temp), although the measured laser temp drops to ~43.5C when the power goes down.

  8299   Fri Mar 15 02:14:27 2013 JenneUpdateCDSSimulink linking to wrong library part

Jamie and I discovered a problem with Matlab/Simulink earlier today. 

In the end suspension models, there is a subblock (with top_names) for ALS stuff.  Inside there, we use a library part called "ALS_END".  When the model was created, it included the part ...../userapps/release/isc/c1/models/ALS_END.mdl .  However, if you open up the c1scy diagram and look in the ALS block for this part, you see the part that is in ..../userapps/release/isc/common/models/ALS_END.mdl .  Note the difference - the one we want is in the c1 directory, while the one that was created (by Jamie) for the LHO One Arm Test is in the common directory. 

If you compile the c1scy model, the RCG is using the correct library part, so the information regarding which part we want is still in there. 

However, if you delete the ALS_END part from the model, put the correct one in, save, close, then reopen the model, it once again displays the wrong model.  The right click "go to library part" option brings you to the library part that is displayed, which is currently the wrong one.  THIS IS BAD, since we could start modifying the wrong things.  You do get a warning by Matlab about the file being "shadowed", so we should take heed when we see that warning, and make sure we are getting the file we want.

We are currently running Matlab version 7.11.0.584, which is r2010b.  Step 1 will be to update Matlab to the latest version, in hopes that this fixes things.  We also should change the name of our c1 part, so that it does not have the same name as the one for the sites.  This is not a great solution since we can't guarantee that we will never choose the same names as other sites, but it will at least fix this one case.  Again, if you see the warning about "shadowed" filenames, pay attention.

  8298   Fri Mar 15 01:57:02 2013 JenneUpdateGreen LockingX arm green locked in TEM00

This work earlier today had required moving the harmonic separator back closer to its original position, so that the green could get through without clipping.  I locked the Xarm (overriding the trigger) and realigned TRX to the PD and camera.

  8297   Thu Mar 14 20:22:33 2013 JenneUpdateGreen LockingXbeat attempt

I aligned the Xgreen and PSL green to overlap on the X beat PD, and reconnected the splitter which combines the X and Y beat signals and sends them to the control room.

I've been stepping the Xend laser temperature offset in steps of 20 counts, making sure the cavity unlocks and relocks on TEM00.  So far I have not seen any beat signals for the Xarm.  I've gone from 0 to 840.

I'll be back in a few hours to keep trying, although interested parties are invited to give it a whirl. 

 

  8296   Thu Mar 14 17:31:57 2013 ManasaUpdateGreen LockingX arm green locked in TEM00

Quote:

Manasa has done some work to get the Xgreen aligned, so I'll switch to trying to find that beatnote for now.

[Jenne, Manasa]

Aligned X-arm green in TEM00.

It was difficult to get the X-arm to lock in TEM00 earlier. Even when it locked, it was just a TEM00 flash. The green was mainly bad in pitch.

I started aligning with the arms flashing in IR and it was still not possible to lock. The second trial was with the arms locked in IR; I lost the green lock when the arms were aligned for IR. I aligned by overlapping the ingoing green with the reflected green visible on the steering mirror. This got the green to lock in TEM00; but still it would only stay that way for 30 sec. Jenne pointed me to Yuta's suggestion of increasing the green refl PD gain. Once I increased the PD gain, the mode stabilized.

I went ahead and centered ALS TRX on the PD and GTRX camera (on the PSL table). ALS_TRX reads 2000+ counts.

EDIT by Jenne:  This corresponds to a power of 550uW on the PSL table, measured before the first out-of-vac steering mirror.

  8295   Thu Mar 14 16:56:16 2013 ChloeUpdate QPD Circuit Design

I have sketched out the circuit design for the QPD. However, it seems like even when using a different opamp configuration, which I talked to Eric about on Friday, space will be a problem. It may be possible to squeeze everything onto a single circuit board to fit in the QPD box but what I think is more likely is that I will need to have 2 separate circuit boards both mounted within the box, one which integrates the signal from the QPD and the other which adds/subtracts (this involves many resistors which will take up a lot of space). I will continue to think about the best design for this.

I will try to have the circuit built in the next week or so, which may be difficult since I just started finals which will take most of my time. I spent most of this week writing up an ECDL proposal for a SURF with Tara. I'll make up for whatever work I miss since I'll be here for my spring break and doing little besides working in lab.

  8294   Thu Mar 14 16:41:48 2013 JenneUpdateGreen LockingYarm ALS laser is funny / dying

Quote (elog 7002, 23July2012):

Jamie and I were doing some locking, and we found that the Yarm green wasn't locking.  It would flash, but not really stay locked for more than a few seconds, and sometimes the green light would totally disappear.  If the end shutter is open, you can always see some green light on the arm transmission cameras.  So if the shutter is open but there is nothing on the camera, that means something is wrong.

I went down to the end, and indeed, sometimes the green light completely disappears from the end table.  At those times, the LED on the front of the laser goes off, then it comes back on, and the green light is back.  This also corresponds to the POWER display on the lcd on the laser driver going to ~0 (usually it reads ~680mW, but then it goes to ~40mW).  The laser stays off for 1-2 seconds, then comes back and stays on for 1-2 minutes, before turning off for a few seconds again.

Koji suggested turning the laser off for an hour or so to see if letting it cool down helps (I just turned it off ~10min ago), otherwise we may have to ship it somewhere for repairs :( 

 This is happening again to the Yend laser.  It's been fine for the afternoon, and I've been playing with the temperature.  First I have been making big sweeps, to figure out what offset values do to the actual temperature, and more recently was starting to do a finer sweep.  Using the 'max hold' function on the 8591, I have seen the beat appear during my big sweeps.  Currently, the laser temperature measurement is at the Yend, and the RF analyzer is here in the control room, so I don't know what temp it was at when the peaks appeared.

Anyhow, while trying to reaquire lock of the TEM00 mode after changing the temperature, I find that it is very difficult (the green seems misaligned in pitch), and every minute or so the light disappears, and I can no longer see the straight-through beam on the camera.  I went down to the end, and the same symptoms of LED on the laser head turning off, power out display goes to ~40mW, are happening.  I have turned off the laser as was the solution last time, in hopes that that will fix things.

Manasa has done some work to get the Xgreen aligned, so I'll switch to trying to find that beatnote for now.

  8293   Thu Mar 14 13:38:29 2013 JenneUpdateGreen LockingPSL green cleaned up

I ran a cable to the GTRX camera.  It is now input #2.  The videoswitch script input naming is modified to match this:  Input 2 used to be "IFOPO", and is now "GTRX".  Input 28 used to be "GRNT", and is now "GTRY".  Both green trans cameras are available from the video screen.

  8292   Thu Mar 14 11:51:14 2013 JamieUpdateGeneralBeatbox upgraded with output whitening, reinstalled

Quote:

I pulled the beatbox from the 1X2 rack so that I could try to hack in some output whitening filters.  These are shamefully absent because of my mis-manufacturing of the power on the board.

Right now we're just using the MON output.  The MON output buffer (U10) is the only chip in the output section that's stuffed:

2013-03-06-195232_1060x927_scrot.png

The power problem is that all the AD829s were drawn with their power lines reversed.  We fixed this by flipping the +15 and -15 power planes and not stuffing the differential output drivers (AD8672).

It's possible to hack in some resistors/capacitors around U10 to get us some filtering there.  It's also possible to just stuff U9, which is where the whitening is supposed to be, then just jump it's output over to the MON output jack.  That might be the cleanest solution, with the least amount of hacking on the board.

I modified the beatbox according to this plan.  I stuffed the whitening filter stage (U9) as indicated in the schematic (I left out the C26 compensation cap which, according to the AD829 datasheet, is not actually needed for our application).  I also didn't have any 301 ohm resistors so I stuffed R18 with 332 ohm, which I think should be fine.

Instead of messing with the working monitor output that we have in place, I stuffed the J5 SMA connector and wired U9 output to it in a single-ended fashion (ie. I grounded the shield pins of J5 to the board since we're not driving it differentially).  I then connected J5 to the I/Q MON outputs on the front panel.  If there's a problem we can just rewire those back to the J4 MON outputs and recover exactly where we were last week.

It all checks out: 0 dB of gain at DC, 1 Hz zero, 10 Hz pole, with 20 dB of gain at high frequencies.

I installed it back in the rack, and reconnected X/Y ARM ALS beatnote inputs and the delay lines.  The I/Q outputs are now connected directly to the DAQ without going through any SR560s (so we recover four SR560s). 

  8291   Thu Mar 14 04:20:54 2013 JenneUpdateGreen LockingYbeat attempt

I dedicated my evening to trying to get the Ygreen beatnote (the idea being to then get the Xgreen beatnote).

First up was tweaking up the green alignment.  Per Yuta's suggestion, elog 8283, I increased the refl PD gain by 2 clicks (20dB) to keep the lock super stable while improving the alignment.  After I finished, I turned it back to its nominal value.  I discovered that I need lenses in front of the DC PD (for Ygreen, and I'm sure Xgreen will be the same).  The beam is just barely taking up the whole 2mm diode, so beam jitter translates directly to DC power change measured by the diode.  I ended up going just by the green transmission camera for the night, and achieved 225uW of Ygreen on the PSL table.  This was ~2,000 counts, but some of the beam is always falling off the diode, so my actual counts value should be higher after installing a lens. 

I then opened up the PSL green shutter, which is controlled by the button labeled "SPS" on the shutter screen - I will fix the label during some coffee break tomorrow.  Using my convenient new PSL green setup, removing the DC PD allows the beam to reflect all the way to the fuse box on the wall, so you can check beam overlap between the PSL green and the arm green at a range of distances.  I did this for Ygreen, and overlapped the Ygreen and PSL green. 

I checked the situation of the beat cabling, since Jamie has the beatbox out for whitening filter modifications tonight.  In order to get some signal into the control room, I connected the output of the BBPD amplifier (mounted on the front of the 1X2 rack) directly to the cable that goes to the control room.  (As part of my cleanup, I put all the cables back the way I found them, so that Jamie can hook everything back up like normal when he finishes the beatbox.) 

I then started watching the signal on the 8591E analyzer, but didn't magically see a peak (one can always hope....).

I decided that I should put the offset in the Y AUX laser slow servo back to the value that we had been using for a long time, ~29,000 counts.  This is where things started going south.  After letting that go for a minute or two, I thought to go check the actual temperature of the laser head.  The "T+" temperature on the controller read something like 42C, but the voltmeter which reads a voltage proportional to the temp (10C/V) was reading 5.6V.  I immediately turned off the offset, but it's going to take a while for it to cool down, so I'll come back in the morning.  I want the AUX laser to be something like 34C, so I just have to wait.  Ooops.

Still to do (for the short-term FPMI):

* Find Y beatnote.

* Align Xgreen to the arm - it's still off in pitch.

* Align Xgreen and PSL green to be overlapped, hitting the BBPD.

* Find the X beatnote.

* Reinstall the beatbox.

* Use ALS to stabilize both arms' lengths.

* Lock MICH with AS.

* Look at the noise spectrum of AS - is there more noise than we expect (Yuta and Koji saw extra noise last summer), and if so, where does it come from?  Yuta calculated (elog 6931) that the noise is much more than expected from just residual arm motion.

* Write a talk.

ELOG V3.1.3-